OpenStudy (anonymous):

log33 + log327 log381 = 4 log330 = 10 log381 = 27 log330 = 3.1

ganeshie8 (ganeshie8):

use below\[\large \log_a m + \log_a n = \log_a (m\times n)\]

ganeshie8 (ganeshie8):

\[\large \log_3 3 + \log_3 27 = ?\]

OpenStudy (anonymous):

log330=10

ganeshie8 (ganeshie8):

nope, its easy actually.. you just need to memorize 2-3 rules and just them : \[\large \log_3 3 + \log_3 27 = \log_3 (3\times 27) \]

ganeshie8 (ganeshie8):

simplifying you get : \[\large \log_3 3 + \log_3 27 = \log_3 (3\times 27) = \log_3 (81)\]

ganeshie8 (ganeshie8):

it simplifies further but you need to use two other rules : \[\large \log_a m^n = n\log _a m\] and \[\large \log_a a = 1\]

OpenStudy (anonymous):

so log381=4

OpenStudy (anonymous):

I made a logarithm tutorial not so long ago, I think you should definitely give it a read, it's pretty friendly :) http://openstudy.com/users/iambatman#/updates/5403fdfde4b0f2ed1e14206a

OpenStudy (anonymous):

am i right ? and i will check it out